OpenStudy (anonymous):

The vertices of the ellipse given by x2+4y2-2x=0 are

OpenStudy (anonymous):

use the formula !

OpenStudy (anonymous):

which formula

OpenStudy (anonymous):

ok first of all you have to show the given equation into standar form !

OpenStudy (anonymous):

like \[x^2/a^2+y^2/b^2 =1\]

OpenStudy (anonymous):

or \[(x-h)^2/a^2 + (y-k)^2/b^2 =1 \]

OpenStudy (anonymous):

(x-1) ^2/a^2+(2y-0)^2/b^2 = 1

OpenStudy (anonymous):

can you show me how u got that equation !

OpenStudy (anonymous):

x^2-2x+(2y) ^2=0 x^2-2x+1+(2y)^2=1

OpenStudy (anonymous):

(x-1) ^ 2+(2y-0)^2=1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then how u get a^2 and b^2

OpenStudy (anonymous):

The base formula for ellipse is \[\frac{ (x-h) ^{2} }{ a^{2} } + \frac{ (y-k) ^{2} }{ b^{2} } = 1\] Vertices of ellipse is f Let see the given equation : \[x^{2}+4y^{2}-2x = 0\] It is similar to : \[x^{2}-2x+4y^{2}=0\] Let change it to: \[x^{2}-2x+1-1+4y^{2}=0\] Change to right form: \[(x^{2}-2x+1) - 1 + 4y^{2}=0\] It is similar to: \[(x-1)^{2} - 1 + 4y^{2}=0\] then: \[(x-1)^{2} + 4y^{2}=1\] Change it to be more closer to base ellipse formula: \[\frac{ (x-1)^{2} }{ 1 } + \frac{ (y)^{2} }{ \frac{ 1 }{ 4 } } = 1\] from the last form we got that the center (h,k) = (1,1) the value of \[a^{2} = 1\], then a = \[\pm 1\] The value of a also tells that the vertices are 1 units to either side of the center, at (h-a, k) and (h+a, k) which are (0,1) and (2,1)